OpenStudy (anonymous):

Need help, I told my friend I would do his homework but turns out I don't know his answer. please help. Thanks

OpenStudy (anonymous):

Ryan, Jessica, and Mark each solved this system of equations. All of their answers are different. Find and explain the mistakes. What is the correct answer? y = 3x + 2 6x – 2y = 10

OpenStudy (anonymous):

Ryan’s answer: I solved by adding the equations. The solution is . (7/6, -3/2). y = 3x + 2 → 3x + y = 2 6x – 2y = 10 → 3x – y = 5 6x = 7 x = 7/6 6x – 2y = 10 → 6(7/6) – 2y = 10 7 – 2y = 10 –2y = –3 y = - 3/2

OpenStudy (anonymous):

Jesse’s answer: I used matrices. The solution is (7/6,7/2) system determinant (3 6 -1 -2)=6-(-6)=12 x determinant (-2 10 -1 -2)=4(-10)=14 y determinant (3 6 -2 10)=30-(-12)=42 x=14/12=7/6 y=42/12=7/2

OpenStudy (anonymous):

I found out marks answer but need jesse's and ryans. thnkx

OpenStudy (mathmale):

Was this meant to be a joke? You are doing someone else's homework and asking others for help in doing so? Please encourage your friend to join OpenStudy and to do his or her own work. Seriously. mathmale, Moderator

OpenStudy (denisaboichuk):

y=42/12=7/2